if a wheel of 1.2m in diameter rotates at one revolution per seconds,calculate the velocity of the wheel?
1
Expert's answer
2018-03-12T16:39:08-0400
The velocity of the wheel
v=ωR
ω=2πn
R=d/2
So
v=2πn d/2=πdn=3.14×1.2×1=3.77 m/s